This is my algebra 2 question. It says to find the value of x. Area of rectangle=36cm2. Xcm and (x+5) how do u get the answer 4.
Answer by ankor@dixie-net.com(22740)

his is my algebra 2 question. It says to find the value of x.
Area of rectangle=36cm2. Xcm and (x+5) how do u get the answer 4.
:
Assume you mean the rectangle is x by (x+5), so you have
x*(x+5) = 36
x^2 + 5x - 36 = 0
Factors to
(x+9)(x-4) = 0
the positive solution
x = 4